# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ E.1.1 Setting $\tau_{\text{min}}$ and extending to models with unknown training data
We can be fairly confident that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on at least some books: the Pile [47] and several other common LLM pre-training datasets either directly include Books3 or other book corpora [75, 76]. Indeed, other work on extraction with uncertain training data treats the Pile as part of a proxy for the true training set [93, 94]. However, without disclosure, we cannot be absolutely certain. And without knowing whether Books3 was in fact included, we cannot—without additional validity experiments—make claims about memorization that are as strong as those we make for models known to be trained on Books3. This is because membership in the training data is a prerequisite for memorization; by definition, a model cannot memorize sequences it has never seen.This is also why observing $p_{\bm{z}}$ on non-training data is important for setting $\tau_{\text{min}}$. By comparing suffix probabilities from negative controls on non-training data (Section 5.2, Appendix E.4) with those from baselines on known training data (Section 5.1, Appendix E.3), we can draw a conservative line for what $p_{\bm{z}}$ is high enough to be indicative of memorization.That is, our negative controls clarify how to set the threshold that we use to count extraction success, $\tau_{\text{min}}$—the minimum $p_{\bm{z}}$ at which we interpret our measurements as evidence of memorization. Running our procedure on non-training data, we confirm that the values of $p_{\bm{z}}$ we observe are far below our chosen cutoff. That way, when we count extraction success for a given sequence with $p_{\bm{z}}\geq\tau_{\text{min}}$, we can be confident that it does not reflect chance generation. Our results in Section 5 and the rest of this Appendix confirm this: $\tau_{\text{min}}=0.1\%$ is highly conservative——well above any $p_{\bm{z}}$ observed in negative control experiments.We therefore use these results to justify our choice of $\tau_{\text{min}}$ in this paper. More extensive experiments would be needed to lower $\tau_{\text{min}}$ or calibrate it model by model. Here, we prefer to be conservative: we would rather under-count memorization than risk over-counting, and defer fine-grained threshold settings to future work.
@@ E.1.2 Relating greedy-decoded and probabilistic discoverable extraction
To our knowledge, we run the largest set of validity experiments in the ML literature on memorization and extraction. Such experiments are not typically performed in work on greedy-decoded discoverable extraction [93, 94, 16]. Hayes et al. [60] conducted one small-scale test on non-training data for a small Pythia model, and used this to support validity claims for probabilistic extraction from much larger and higher quality Llama 1 models. Given these norms, we initially omitted our negative controls from the main paper, deferring discussion of Phi 4 experiments to the Appendix and omitting our varied-prefix experiments entirely from the paper. However, after receiving some feedback about earlier drafts of this work that indicated serious misunderstandings about the nature of memorization and extraction, we realized that including the results of these experiments in the main text could help some readers better understand our results and claims. We now clearly delineate two sets of negative controls, both in this appendix and in the main paper text (Section 5.2): the experiments on Phi 4 (Appendix E.4.1) and the experiments on books published after training-date cutoff (Appendix E.4.2). Here, we clarify the relationship between prior definitions and our approach, and explain how this motivates our validity experiments.
@@
It is widely accepted that successful greedy-decoded discoverable extraction on known training data is evidence of memorization; memorization remains the most parsimonious explanation [60, 16, 93, 15, 28]. In this procedure, greedy decoding always selects the top-ranked token at each step, making it a deterministic approximation of the most likely sequence. Carlini et al. [16] describe this choice as a tractable proxy for the $\operatorname*{arg\,max}$ in their definition of model knowledge extraction:
Definition 3 (Model Knowledge Extraction, from Carlini et al. [16]).
@@
We add emphasis to two areas in this definition, which we discuss below. First, we note that this definition does not mention training data; it is just a definition about a model having knowledge of string $s$. Carlini et al. [16] connect extraction of training data to memorization through a second definition:
Definition 4 ($k$-Eidetic Memorization, from Carlini et al. [16]).
@@
Greedy decoding is equivalent to top-$k$ decoding with T=1T{=}1 and k=1k={1}; one uses the base model distribution (T=1T{=}1) and then only considers the token with top-$1$-ranked logit (k=1k{=}1). Probabilistic extraction relaxes this by considering a wider set of high-probability continuations (e.g., top-$40$). This better reflects the model’s distribution and avoids the problem where greedy decoding yields globally low-probability sequences. Hayes et al. [60] motivate their metric precisely from this point: the greedy-decoded suffix is locally optimal, but it may lead to a globally relatively low probability sequence under the model.A hypothetical example illustrates this gap: a greedy-decoded suffix can have vanishing total probability, while a slightly different path within the top-$k$ yields a much higher-probability suffix. For this example, probabilities are computed with respect to the base $T=1$ distribution. Consider a greedy-decoded suffix where the first token has $99.9\%$ probability and is selected, and then every top-$1$ token after it for $49$ tokens has probability $1\%$; the total probability for the $50$-token suffix would be $999\times(0.01)^{49}{\approx}9.99\times 10^{-99}.$ Now imagine we instead were not restricted to always picking the top-$1$ probability token; we relax this requirement to, for example, allow to select from the top-$k$ tokens. Again we consider probabilities computed with respect to the base $T=1$ distribution, so that sequence probabilities are computed off the same logits. For the first token position, the top-$2$ token might have probability $1\%$, and then the remaining top-$1$ $49$ tokens could each have probability $99.9\%$; the total probability for this suffix would be $01\times(0.999)^{49}{\approx}0.001$ (or about $0.1\%$). The point is, by allowing for a bit more flexibility in the decoding strategy, the model can potentially produce much higher probability sequences; when we observe that such (high-enough-probability, $\geq\tau_{\text{min}}$) sequences match the verbatim suffix from the training data, then we count them as extracted. As long as we choose $\tau_{\text{min}}$ appropriately—i.e., such that it does not capture low-probability sequences that are perhaps reflective of generalization—we should obtain higher fidelity measurements of memorization. We should have fewer false negatives than greedy-decoded discoverable extraction. And notably, as we show in Appendix A.3, this is actually cheaper to compute than how work like Carlini et al. [16] computes extraction, as we can compute probabilistic extraction directly off of the logits, rather than needing to actually generate completions.
@@ E.1.3 The role of prefix length in our experiments
We run our validity experiments with various prefix lengths, not just for $100$-token sequences with $50$-token prefixes and $50$-token suffixes. This is a deliberate choice because additional context biases the model’s output distribution toward particular continuations. For sequences from the training data, this is known to surface memorization that may not appear for shorter prefixes. Carlini et al. [18] call this the discoverability phenomenon: “some memorization only becomes apparent …when the model is prompted with a sufficiently long context.” Through the lens of probabilistic extraction, a longer training-data prefix can raise the probability of the target suffix in Equation 1, such that it surpasses $\tau_{\text{min}}$. (It is worth noting that any conditional extraction procedure—i.e., one that uses prompts, rather than unconditionally generating from the LLM—statistically biases the output. This is one of the reasons that extraction under-counts memorization; it only searches a part of the learned distribution.)By testing various prefix lengths on known training data, we can estimate an upper bound on memorization (with respect to our extraction procedure) for a given work. We expect longer prefixes to expose more memorization, but that increasing prefix length will not typically induce all suffix probabilities to $100\%$; that is, we will observe more extraction only if there is more memorization to uncover. It is not the case that one can simply make the prefix arbitrarily long and then extract any target suffix. We discuss this further in Appendices E.3 and E.5. This logic also strengthens our negative controls. In this case, we should not be able to measure extraction for any unique sequences of non-training data—even when we intentionally induce more bias with longer prompts. Again, this shows that arbitrarily long suffixes do not all of a sudden mean we will register extraction with our measurements. Here, we should not be able to do so because the sequences were not included in the training data.
@@ E.2 Setup
Running experiments with longer sequences is significantly more expensive. We therefore limit our results on varied prefixes to a small selection of books from Books3: Beloved [89], Harry Potter and the Sorcerer’s Stone [106], This Is How You Lose Her [40], and We Were Eight Years in Power [23]. We picked these $4$ books because each represents (roughly speaking) a different pattern for memorization in our sliding-window results on $50$-token prefixes (Appendix D) for the model that exhibits the most memorization: Llama 3.1 70B. This Is How You Lose Her is barely memorized at all (Appendix D.3.12. Beloved has fragmented memorization “hot-spots;” there are many throughout the book, but they are rather disconnected, separated by lots of regions that do not register any memorization (Appendix D.3.34). We Were Eight Years in Power has several “hot-spots;” some are fragmented, but some indicate long, continuous regions of memorized text (Appendix D.3.7. Harry Potter and the Sorcerer’s Stone is effectively memorized entirely, though not all suffixes have equally high extraction probability (Appendix D.3.39).
@@
1 Gareth Brown The Society of Unknowable Objects [7] July 31, 2025 ©
2 Emily Henry Great Big Beautiful Life [63] April 22, 2025 ©
3 Ocean Vuong The Emperor of Gladness [130] May 13, 2025 ©
4 Sarah Wynn-Williams Careless People [136] March 11, 2025 ©
@@
Following from above, we run these experiments on $3$ models known to have been trained on Books3. We include Llama 3.1 70B, because this is the model that has memorized the most of all Llama models we test. We also include Llama 3.1 8B, as it is a much smaller model from the same family. It memorizes less than Llama 3.1 70B. (These results omitted for brevity in Appendix D.) While it may be possible that increasing prefix length for Llama 3.1 70B surfaces many more books that are effectively entirely memorized (i.e., like Harry Potter and the Sorcerer’s Stone), we do not expect this to be the case for Llama 3.1 8B, since it is a much smaller capacity model. Together, this will clarify how it is not the case that just making prefixes longer means that one could generate any target suffix they so desire—regardless of whether it was memorized. We also include Llama 2 13B. This model is a similar size as Phi 4, which we use in our negative controls. They have similar capacity, but were clearly trained very differently. While we do not know all of the differences, we do know that Llama 2 13B was trained on Books3, while Phi 4 was not. For one experiment on non-training data, we also include Qwen 2.5 72B as a point of comparison.
@@ E.3 Results of varying prefix length on known training data
We next provide extended results for baseline experiments that vary prefix lengths on known training data. Our results yield the outcomes we expect: we are able to surface more memorization with increased prefix length—to a point. Except in rare cases like Harry Potter and the Sorcerer’s Stone and Llama 3.1 70B, in which the entire book is memorized, arbitrarily increasing prefix length does not make very training data suffix extractable. This makes sense, as not all training data are memorized; our measurements should therefore not register all training sequences as extractable. These results also make clear that absence of extraction is not itself evidence that a sequence is not a member of the training data. Most sequences from Books3 are not extractable even from models like Llama 3.1 70B, which memorize a lot. So the claim we make with respect to absence of extraction signal is in this one direction. When we make claims about training data membership, these are either based on ground-truth knowledge (e.g., knowing the Books3 was in the training data and that 2025 books were not), or conclusions that we draw based on comparisons to our negative controls (Appendix E.6).We include results for $3$ different models on $4$ books from Books3, as documented in Appendix E.2. The largest model we test is Llama 3.1 70B (Figure 115). For a point of comparison with a much smaller capacity model, we also include the smaller model from the same family—Llama 3.1 8B (Figure 116) and Llama 2 13B (Figure 117), since this model is of a similar size as Phi 4 (Section 5.2, Appendix E.4.1).
@@
Figure 115: Llama 3.1 70B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [89], Harry Potter and the Sorcerer’s Stone [106], This Is How You Lose Her [40], and We Were Eight Years in Power [23]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 3.1 70B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.
Figure 116: Llama 3.1 8B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [89], Harry Potter and the Sorcerer’s Stone [106], This Is How You Lose Her [40], and We Were Eight Years in Power [23]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 3.1 8B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.
@@
Figure 117: Llama 2 13B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [89], Harry Potter and the Sorcerer’s Stone [106], This Is How You Lose Her [40], and We Were Eight Years in Power [23]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 2 13B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.

### v4

@@ E.1.1 Setting $\tau_{\text{min}}$ and extending to models with unknown training data
We can be fairly confident that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on at least some books: the Pile [gao2020pile] and several other common LLM pre-training datasets either directly include Books3 or other book corpora [lee2023explainers, lee2023talkin]. Indeed, other work on extraction with uncertain training data treats the Pile as part of a proxy for the true training set [nasr2023scalable, nasr2025scalable]. However, without disclosure, we cannot be absolutely certain. And without knowing whether Books3 was in fact included, we cannot—without additional validity experiments—make claims about memorization that are as strong as those we make for models known to be trained on Books3. This is because membership in the training data is a prerequisite for memorization; by definition, a model cannot memorize sequences it has never seen.
This is also why observing $p_{\bm{z}}$ on non-training data is important for setting $\tau_{\text{min}}$. By comparing suffix probabilities from negative controls on non-training data (Section 5.2, Appendix E.4) with those from baselines on known training data (Section 5.1, Appendix E.3), we can draw a conservative line for what $p_{\bm{z}}$ is high enough to be indicative of memorization.
That is, our negative controls clarify how to set the threshold that we use to count extraction success, $\tau_{\text{min}}$—the minimum $p_{\bm{z}}$ at which we interpret our measurements as evidence of memorization. Running our procedure on non-training data, we confirm that the values of $p_{\bm{z}}$ we observe are far below our chosen cutoff. That way, when we count extraction success for a given sequence with $p_{\bm{z}}\geq\tau_{\text{min}}$, we can be confident that it does not reflect chance generation. Our results in Section 5 and the rest of this Appendix confirm this: $\tau_{\text{min}}=0.1\%$ is highly conservative——well above any $p_{\bm{z}}$ observed in negative control experiments.
We therefore use these results to justify our choice of $\tau_{\text{min}}$ in this paper. More extensive experiments would be needed to lower $\tau_{\text{min}}$ or calibrate it model by model. Here, we prefer to be conservative: we would rather under-count memorization than risk over-counting, and defer fine-grained threshold settings to future work.
@@ E.1.2 Relating greedy-decoded and probabilistic discoverable extraction
To our knowledge, we run the largest set of validity experiments in the ML literature on memorization and extraction. Such experiments are not typically performed in work on greedy-decoded discoverable extraction [nasr2023scalable, nasr2025scalable, carlini2021extracting]. hayes2025measuringmemorizationlanguagemodels conducted one small-scale test on non-training data for a small Pythia model, and used this to support validity claims for probabilistic extraction from much larger and higher quality Llama 1 models.
Given these norms, we initially omitted our negative controls from the main paper, deferring discussion of Phi 4 experiments to the Appendix and omitting our varied-prefix experiments entirely from the paper. However, after receiving some feedback about earlier drafts of this work that indicated serious misunderstandings about the nature of memorization and extraction, we realized that including the results of these experiments in the main text could help some readers better understand our results and claims. We now clearly delineate two sets of negative controls, both in this appendix and in the main paper text (Section 5.2): the experiments on Phi 4 (Appendix E.4.1) and the experiments on books published after training-date cutoff (Appendix E.4.2).
Here, we clarify the relationship between prior definitions and our approach, and explain how this motivates our validity experiments.
@@
It is widely accepted that successful greedy-decoded discoverable extraction on known training data is evidence of memorization; memorization remains the most parsimonious explanation [hayes2025measuringmemorizationlanguagemodels, carlini2021extracting, nasr2023scalable, carlini2025blog, cooper2024files]. In this procedure, greedy decoding always selects the top-ranked token at each step, making it a deterministic approximation of the most likely sequence. carlini2021extracting describe this choice as a tractable proxy for the $\operatorname*{arg\,max}$ in their definition of model knowledge extraction:
Definition 3 (Model Knowledge Extraction, from carlini2021extracting).
@@
We add emphasis to two areas in this definition, which we discuss below. First, we note that this definition does not mention training data; it is just a definition about a model having knowledge of string $s$. carlini2021extracting connect extraction of training data to memorization through a second definition:
Definition 4 ($k$-Eidetic Memorization, from carlini2021extracting).
@@
Greedy decoding is equivalent to top-$k$ decoding with T=1T{=}1 and k=1k={1}; one uses the base model distribution (T=1T{=}1) and then only considers the token with top-$1$-ranked logit (k=1k{=}1). Probabilistic extraction relaxes this by considering a wider set of high-probability continuations (e.g., top-$40$). This better reflects the model’s distribution and avoids the problem where greedy decoding yields globally low-probability sequences. hayes2025measuringmemorizationlanguagemodels motivate their metric precisely from this point: the greedy-decoded suffix is locally optimal, but it may lead to a globally relatively low probability sequence under the model.
A hypothetical example illustrates this gap: a greedy-decoded suffix can have vanishing total probability, while a slightly different path within the top-$k$ yields a much higher-probability suffix. For this example, probabilities are computed with respect to the base $T=1$ distribution. Consider a greedy-decoded suffix where the first token has $99.9\%$ probability and is selected, and then every top-$1$ token after it for $49$ tokens has probability $1\%$; the total probability for the $50$-token suffix would be $999\times(0.01)^{49}{\approx}9.99\times 10^{-99}.$ Now imagine we instead were not restricted to always picking the top-$1$ probability token; we relax this requirement to, for example, allow to select from the top-$k$ tokens. Again we consider probabilities computed with respect to the base $T=1$ distribution, so that sequence probabilities are computed off the same logits. For the first token position, the top-$2$ token might have probability $1\%$, and then the remaining top-$1$ $49$ tokens could each have probability $99.9\%$; the total probability for this suffix would be $01\times(0.999)^{49}{\approx}0.001$ (or about $0.1\%$).
The point is, by allowing for a bit more flexibility in the decoding strategy, the model can potentially produce much higher probability sequences; when we observe that such (high-enough-probability, $\geq\tau_{\text{min}}$) sequences match the verbatim suffix from the training data, then we count them as extracted. As long as we choose $\tau_{\text{min}}$ appropriately—i.e., such that it does not capture low-probability sequences that are perhaps reflective of generalization—we should obtain higher fidelity measurements of memorization. We should have fewer false negatives than greedy-decoded discoverable extraction. And notably, as we show in Appendix A.3, this is actually cheaper to compute than how work like carlini2021extracting computes extraction, as we can compute probabilistic extraction directly off of the logits, rather than needing to actually generate completions.
@@ E.1.3 The role of prefix length in our experiments
We run our validity experiments with various prefix lengths, not just for $100$-token sequences with $50$-token prefixes and $50$-token suffixes. This is a deliberate choice because additional context biases the model’s output distribution toward particular continuations. For sequences from the training data, this is known to surface memorization that may not appear for shorter prefixes. carlini2023quantifying call this the discoverability phenomenon: “some memorization only becomes apparent …when the model is prompted with a sufficiently long context.” Through the lens of probabilistic extraction, a longer training-data prefix can raise the probability of the target suffix in Equation 1, such that it surpasses $\tau_{\text{min}}$. (It is worth noting that any conditional extraction procedure—i.e., one that uses prompts, rather than unconditionally generating from the LLM—statistically biases the output. This is one of the reasons that extraction under-counts memorization; it only searches a part of the learned distribution.)
By testing various prefix lengths on known training data, we can estimate an upper bound on memorization (with respect to our extraction procedure) for a given work. We expect longer prefixes to expose more memorization, but that increasing prefix length will not typically induce all suffix probabilities to $100\%$; that is, we will observe more extraction only if there is more memorization to uncover. It is not the case that one can simply make the prefix arbitrarily long and then extract any target suffix. We discuss this further in Appendices E.3 and E.5.
This logic also strengthens our negative controls. In this case, we should not be able to measure extraction for any unique sequences of non-training data—even when we intentionally induce more bias with longer prompts. Again, this shows that arbitrarily long suffixes do not all of a sudden mean we will register extraction with our measurements. Here, we should not be able to do so because the sequences were not included in the training data.
@@ E.2 Setup
Running experiments with longer sequences is significantly more expensive. We therefore limit our results on varied prefixes to a small selection of books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. We picked these $4$ books because each represents (roughly speaking) a different pattern for memorization in our sliding-window results on $50$-token prefixes (Appendix D) for the model that exhibits the most memorization: Llama 3.1 70B. This Is How You Lose Her is barely memorized at all (Appendix D.3.12. Beloved has fragmented memorization “hot-spots;” there are many throughout the book, but they are rather disconnected, separated by lots of regions that do not register any memorization (Appendix D.3.34). We Were Eight Years in Power has several “hot-spots;” some are fragmented, but some indicate long, continuous regions of memorized text (Appendix D.3.7. Harry Potter and the Sorcerer’s Stone is effectively memorized entirely, though not all suffixes have equally high extraction probability (Appendix D.3.39).
@@
1 Gareth Brown The Society of Unknowable Objects [dante] July 31, 2025 ©
2 Emily Henry Great Big Beautiful Life [greatbig] April 22, 2025 ©
3 Ocean Vuong The Emperor of Gladness [emperorgladness] May 13, 2025 ©
4 Sarah Wynn-Williams Careless People [careless] March 11, 2025 ©
@@
Following from above, we run these experiments on $3$ models known to have been trained on Books3. We include Llama 3.1 70B, because this is the model that has memorized the most of all Llama models we test. We also include Llama 3.1 8B, as it is a much smaller model from the same family. It memorizes less than Llama 3.1 70B. (These results omitted for brevity in Appendix D.) While it may be possible that increasing prefix length for Llama 3.1 70B surfaces many more books that are effectively entirely memorized (i.e., like Harry Potter and the Sorcerer’s Stone), we do not expect this to be the case for Llama 3.1 8B, since it is a much smaller capacity model. Together, this will clarify how it is not the case that just making prefixes longer means that one could generate any target suffix they so desire—regardless of whether it was memorized.
We also include Llama 2 13B. This model is a similar size as Phi 4, which we use in our negative controls. They have similar capacity, but were clearly trained very differently. While we do not know all of the differences, we do know that Llama 2 13B was trained on Books3, while Phi 4 was not. For one experiment on non-training data, we also include Qwen 2.5 72B as a point of comparison.
@@ E.3 Results of varying prefix length on known training data
We next provide extended results for baseline experiments that vary prefix lengths on known training data. Our results yield the outcomes we expect: we are able to surface more memorization with increased prefix length—to a point. Except in rare cases like Harry Potter and the Sorcerer’s Stone and Llama 3.1 70B, in which the entire book is memorized, arbitrarily increasing prefix length does not make very training data suffix extractable. This makes sense, as not all training data are memorized; our measurements should therefore not register all training sequences as extractable. These results also make clear that absence of extraction is not itself evidence that a sequence is not a member of the training data. Most sequences from Books3 are not extractable even from models like Llama 3.1 70B, which memorize a lot. So the claim we make with respect to absence of extraction signal is in this one direction. When we make claims about training data membership, these are either based on ground-truth knowledge (e.g., knowing the Books3 was in the training data and that 2025 books were not), or conclusions that we draw based on comparisons to our negative controls (Appendix E.6).
We include results for $3$ different models on $4$ books from Books3, as documented in Appendix E.2. The largest model we test is Llama 3.1 70B (Figure 115). For a point of comparison with a much smaller capacity model, we also include the smaller model from the same family—Llama 3.1 8B (Figure 116) and Llama 2 13B (Figure 117), since this model is of a similar size as Phi 4 (Section 5.2, Appendix E.4.1).
@@
Figure 115: Llama 3.1 70B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 3.1 70B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.
Figure 116: Llama 3.1 8B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 3.1 8B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.
@@
Figure 117: Llama 2 13B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 2 13B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.